I just need the correct format on how to solve this problem.. I tend to guess and and chicken scratch. Just looking for the proper formula. Thanks
The second angle of a triangle is 4 times as large as the first. The third angle is 80 degrees more than the sum of the other two angles. Find the measure of the second angle.

Answer by scott8148(6628)

let x = "first angle"

"The second angle of a triangle is 4 times as large as the first" ___ 4 * x = 4x

"The third angle is 80 degrees more than the sum of the other two angles" ___ x + 4x + 80 = 5x + 80

x + 4x + 5x + 80 = 180

10x + 80 = 180 ___ 10x = 100 ___ x = 10

second angle is 4 * 10 or 40